OpenStudy (anonymous):

If a 30N net force acts on a 20kg object for 5s how fast will the object be moving at the time and how far will it have traveled?

OpenStudy (anonymous):

F=M*A 30N=(20kg)*A 30/20= 1.5 Acceleration=1.5ms^-2 then plug that into this equation S=UT+1/2AT^2 S=0(5) + 1/2(1.5)(5)^2 S=18.75meters 18.75 is the distance traveled in 5 seconds now to solve for velocity Vf=Vi+AT Vf=0+(1.5)(5) Vf=7.5m/s^-1 7.5m/s^-1 is the velocity at 5 seconds